OpenStudy (anonymous):

Change (-5, -7) into polar form. Round both r and θ to the nearest tenth. Help plz??

OpenStudy (anonymous):

this is how i did it..but i'm not sure

OpenStudy (anonymous):

x = -5 y = -7 r = sqrt(x^2 + y^2) = sqrt r = sqrt(x^2 + y^2) = sqrt(74) = 8.60 theta = tan^1 (y/x) = tan^1 (5/7) = 35.5 degrees. The polar coordinates are approximately (8.6, 35.5 degrees)

terenzreignz (terenzreignz):

Given the rectangular coordinates, (x, y), To get the polar coordinates (r, θ), use these formulas \[r = \sqrt{x^{2} + y^{2}}\] and (I'm assuming you have a calculator at your disposal) \[\theta=\tan^{-1}\left| \frac{y}{x} \right|\]if BOTH x and y are positive\[\theta=180°-\tan^{-1}\left| \frac{y}{x} \right|\]if x is negative and y is positive\[\theta=180°+\tan^{-1}\left| \frac{y}{x} \right|\]if BOTH x and y are negative and finally\[\theta=360°-\tan^{-1}\left| \frac{y}{x} \right|\]if x is positive and y is negative

terenzreignz (terenzreignz):

You seem to have gotten r down well enough, but remember, given that both x and y are negative, and your angle is less than 90 degrees, that means your coordinates are in the FIRST quadrant, is that really the case?

OpenStudy (anonymous):

ohhh ok

terenzreignz (terenzreignz):

See, both x and y are negative, and your calculator only displays values from 0 to 90 for inverse trigs (I think?) So better use that third formula for theta to ensure that it ends up in the third quadrant as it should :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

so i'll use theta = 180 degrees - tan one or?

OpenStudy (anonymous):

anybody??